# How to Train Long-Context Language Models (Effectively)

### v3

@@ B.5 Synthetic data with a stronger data generator
B.6 Comparison to Fu et al. (2024)
We show a head-to-head comparison to the data strategy of Fu et al. (2024) in Table 24. We see that under a fair comparison, our data mix significantly outperforms Fu et al. (2024) on both short and long-context tasks. The main difference of the two data strategies is that Fu et al. (2024) proportionally up-sample long documents in each domain with an arbitrary ratio; ProLong uses a mix of short and long documents, where the ratio of the mix and the domains for the long documents are carefully ablated.
Table 24: Comparison between Fu et al. (2024) and our model. For a fair comparison, we use the same initialization (Llama-3-8B), same amount of data (5B), and same hyperparameters (§A.4). The ProLong data mix significantly outperforms Fu et al. (2024) on both short and long-context tasks.
@@ B.7 Short-context performance after SFT
Table 25: Short-context performance of our model after SFT. We also report a baseline using Llama-3-8B as the initialization and data from Fu et al. (2024), trained with 5B tokens. ProLong is initialized from Llama-3-8B-Instruct. “Llama-3-8B-Instruct + UltraChat”: for a more fair comparison to ProLong, we conduct SFT on top of Llama-3-8B-Instruct with UltraChat. ProLong largely retraines the short-context performance of Llama-3-8B-Instruct except MMLU and GSM8K. We hypothesize that the close-source instruction tuning data of Llama-3-8B-Instruct is heavily engineered to improve math and knowledge-intensive tasks, which we do not have access to. ProLong achieves comparable results to “Llama-3-8B-Instruct + UltraChat”, which further demonstrates that our data mix effective retains short-context performance.

### v4

@@ B.5 Synthetic data with a stronger data generator
B.6 Comparison to fu2024data
We show a head-to-head comparison to the data strategy of fu2024data in Table 24. We see that under a fair comparison, our data mix significantly outperforms fu2024data on both short and long-context tasks. The main difference of the two data strategies is that fu2024data proportionally up-sample long documents in each domain with an arbitrary ratio; ProLong uses a mix of short and long documents, where the ratio of the mix and the domains for the long documents are carefully ablated.
Table 24: Comparison between fu2024data and our model. For a fair comparison, we use the same initialization (Llama-3-8B), same amount of data (5B), and same hyperparameters (§A.4). The ProLong data mix significantly outperforms fu2024data on both short and long-context tasks.
@@ B.7 Short-context performance after SFT
Table 25: Short-context performance of our model after SFT. We also report a baseline using Llama-3-8B as the initialization and data from fu2024data, trained with 5B tokens. ProLong is initialized from Llama-3-8B-Instruct. “Llama-3-8B-Instruct + UltraChat”: for a more fair comparison to ProLong, we conduct SFT on top of Llama-3-8B-Instruct with UltraChat. ProLong largely retraines the short-context performance of Llama-3-8B-Instruct except MMLU and GSM8K. We hypothesize that the close-source instruction tuning data of Llama-3-8B-Instruct is heavily engineered to improve math and knowledge-intensive tasks, which we do not have access to. ProLong achieves comparable results to “Llama-3-8B-Instruct + UltraChat”, which further demonstrates that our data mix effective retains short-context performance.